17. You are lying 120 ft away from a tree that is 50 feet tall. You look up at the top of the tree. Approximately how far is your hear from the top of the tree in a straight line
Answer by nerdybill(7384)

17. You are lying 120 ft away from a tree that is 50 feet tall. You look up at the top of the tree. Approximately how far is your hear from the top of the tree in a straight line
.
Apply Pythagorean theorem:
Let x = distance from ear to top of tree
then
x^2 = 50^2 + 120^2
x^2 = 2500 + 14400
x^2 = 16900
x = sqrt(16900)
x = 130 feet